# Estimating Heterogeneous Causal Effects of High-Dimensional Treatments:

### v4

@@ 5.3 Group Membership
The right panel of the figure shows the distribution of other respondent characteristics. In general, Group 2 in the two-group analysis and Group 3 in the three-group analysis consist of those who live in ZIP codes with few immigrants and have lower educational achievements. For the three-group analysis, those in Group 2 tend to be Republican, whereas those in Group 1 are more likely to be Democrats. This is consistent with the finding of a larger penalty for entry without legal authorization in Group 2. Group 3 contains a mix of political ideologies, though it has more respondents who identify as Undecided/Independent/Other or not strong Republican than the other two groups.
What respondent characteristics are predictive of the group membership? In addition to the covariate distribution for each group shown in Fig 4, we can also find how important each moderator is in predicting group membership, conditional on all other moderators. We examine how the predicted probabilities of group memberships change across respondents with different characteristics. Specifically, we estimate
$\mathbb{E}\left[\pi_{k}(X_{ij}=x_{1},\bm{X}_{i,-j})-\pi_{k}(X_{ij}=x_{0},\bm{X}_{i,-j})\right]$ (10)
where $x_{0}$ and $x_{1}$ are different values of covariate of interest $X_{ij}$. If $X_{ij}$ is a categorical variable, we set $x_{0}$ to the baseline level and $x_{1}$ to the level indicated on the vertical axis. If $X_{ij}$ is a continuous variable as in the case of the Hispanic prejudice score, then $x_{0}$ and $x_{1}$ represent the 25th and 75th percentile values. The solid arrows represent whether the corresponding 95% Bayesian credible interval covers zero or not. Appendix K shows the effect of changing a moderator on the absolute value of the changes in predicted probabilities of group membership. In some cases, changing a moderator shows a small average change but a larger average of absolute changes.
Figure 5: The impact of moderator values on likelihood of being assigned to groups, for two-group (left two plots) and three-group (right three plots) analysis. Dark arrows indicate that there is a significant effect of the moderator on group membership, i.e. that the corresponding quantity defined in Equation (10) is statistically significant.
Consistent with the earlier findings, Fig 5 shows that those with high Hispanic prejudice scores tend to be part of Group 2 in the two-group analysis and Group 3 in the three-group analysis even after controlling for other moderators. These respondents are also less likely to be members of Group 1 in both analyses. For the three-group analysis, party ID plays a statistically significant role (indicated by dark arrow) with Republicans tending to be part of Group 2 or 3 and more Democrats belonging to Group 1. On average, respondents in Group 1 have higher education in both analyses.
@@
All of the AMIE effects found are quite small, so we do not present those results here. According to the skill-premium theory of Newman and Malhotra (2019), we expect to find an interaction between job and country or education in country, in at least some groups. Unfortunately, our analysis does not find support for this hypothesis.
@@ 5.4 Comparison to an alternative method
$\mathrm{CAMCE}_{j}(l,l^{\prime};\bm{X}_{i})=\sum_{k=1}^{K}\delta_{jk}(l,l^{\prime})\pi_{k}(\bm{X}_{i})$ (11)
@@
Appendix B compares the estimated CAMCE obtained from our method and Robinson and Duch (2024)’s (cjbart) using the same moderators and treatments. Our method discovers a considerable degree of heterogeneity in the CAMCEs whereas cjbart shows limited treatment effect variation for most countries. Under our model, the estimated heterogeneous effects are more strongly associated with predictors than cjbart; for example, our method finds a clear association, on average, between the estimated CAMCE and prejudice or party identification whereas cjbart does not.
@@ 6 Concluding Remarks
Figure A1: Comparison of discovered heterogeneous effects (the conditional average marginal component effects or CAMCEs) between the proposed method and the BART-based method cjbart. In both plots the y-axis corresponds to values estimated, either by our method (right) or by cjbart Robinson and Duch (2024) (left). The plots show the estimated effect of Iraq as compared to the baseline of Germany. In the top figure, the x-axis and color corresponds to the categories of individuals based on the quartile of their Hispanic prejudice score. In the bottom figure, the x-axis and color corresponds to party ID.
@@
Lemma C.1 (Deviance Decomposition for Exponential Family)
@@
Proposition D.1 (Finding maximally heterogeneous groups with moderators)
@@
Lemma D.1 (Entropy of Groups with Respect to Moderators)
@@
Supplementary Material E Propriety of the Structured Sparse Prior
@@
By applying Theorem E.1 and noting that the null spaces of $\bm{A}^{T}\bm{A}$ and $\bm{A}$ are identical, the integral of $\tilde{k}(\tilde{\bm{\beta}})$ is finite if and only if $\bm{Q}\mathcal{B}_{\bm{C}^{\top}}$ is full column rank, where $\bm{Q}^{\top}=[\bm{Q}^{\top}_{1},\cdots,\bm{Q}^{\top}_{G}]$. We demonstrate this fact in two steps. First, there exists a permutation matrix $\bm{P}_{Q}$ such that $\bm{P}_{Q}\bm{Q}$ has a block diagonal structure with $J+1$ diagonal blocks. The first $J$ blocks corresponding to the main terms for each factor $j$ and the last block corresponds to all interaction terms. The null space of each block is spanned by the vector $\bm{1}$ as the corresponding block of $\bm{P}_{Q}\bm{Q}$ is a (transposed) orientated incidence matrix of a fully connected graph. Thus, the null space of $\bm{P}_{Q}\bm{Q}$, and hence $\bm{Q}$, is spanned by the $J+1$ columns of a block diagonal matrix with $\bm{1}$ on each block. Second, consider the linear constraints $\bm{C}^{\top}\bm{\beta}=\bm{0}$. The only vector to satisfy this constraint and lie in the null space of $\bm{Q}$ must be $\bm{0}$ as, for each block, the only vector proportional to $\bm{1}$ and satisfying the corresponding sum-to-zero constraints must be $\bm{0}$. Thus, $\bm{Q}\mathcal{B}_{\bm{C}^{\top}}$ is full column rank and the prior is proper. ∎
Supplementary Material F Derivations for the Basic Model
@@
F.1 Removing the Linear Constraints
@@
Given the similarity of the unconstrained and constrained problems and for notational simplicity, we present all results herein dropping the second “tilde” notation on $\tilde{\bm{T}}_{i}$ and the “tilde” on $\bm{\beta}_{k}$ and note that, once estimated, $\tilde{\bm{\beta}}_{k}$ is projected back into the original space for the reported coefficients, average marginal component effects, etc. The results of Appendix I on approximating $\tilde{\bm{\beta}}_{k}$ as multivariate Gaussian imply that $\bm{\beta}_{k}$ will have a (singular) multivariate Gaussian distribution.
F.2 Expectation Maximization Algorithm
This section considers inference after removing the linear constraints as discussed in the prior subsection. Algorithm A1 summarizes our approach to maximizing Equation (9). Each iteration of our AECM algorithm involves two cycles where the data augmentation scheme enables iterative updating of the treatment effect parameters $\bm{\beta}$ and moderators $\bm{\phi}$. $\bm{\theta}$ collects both sets of parameters.
@@ F.2 Expectation Maximization Algorithm
Initialize Parameters: $\bm{\theta}^{(0)}$, i.e. $\bm{\beta}^{(0)}$ and $\bm{\phi}^{(0)}$; Appendix F.4 provides details.
@@
     1a. $E$-Step: Find the conditional distributions of $\{Z_{i},\omega_{i}\}_{i=1}^{N}$ and $\{\{\tau^{2}_{gk}\}_{g=1}^{G}\}_{k=1}^{K}$ given $\{Y_{i}\}$ and $\bm{\theta}^{(t)}$ (Eq. (A1)). Derive $Q_{\beta}(\bm{\beta},\bm{\theta}^{(t)})$ (Eq. (F.2.1)).
@@
     2a. $E$-Step: Find $p(Z_{i}=k\mid~{}Y_{i},\bm{\beta}^{(t+1)},\bm{\phi}^{(t)})$. Derive $Q_{\phi}(\bm{\phi},\{\bm{\beta}^{(t+1)},\bm{\phi}^{(t)}\})$ (Eq. (F.2.2)).
@@
     3. Stop if $\log p(\bm{\theta}^{(t+1)}|\{Y_{i}\})-\log p(\bm{\theta}^{(t)}|\{Y_{i}\})<\epsilon_{1}$ (Eq. (9)) or $||\bm{\theta}^{(t+1)}-\bm{\theta}^{(t)}||_{\infty}<\epsilon_{2}$.
F.2.1 Updating Treatment Effect Parameters
@@
$p(Y_{i},\omega_{i}\mid Z_{i})\ \propto\ \frac{1}{2}\exp\left\{\left(Y_{i}-\frac{1}{2}\right)\psi_{Z_{i}}(\bm{T}_{i})-\frac{\omega_{i}}{2}\left[\psi_{Z_{i}}(\bm{T}_{i})\right]^{2}\right\}f_{PG}(\omega_{i}\mid 1,0),$ (A1a)
@@
where all expectations are taken over the conditional distribution of the missing data given the current parameter estimates. We note that the $E$-Step involves computing $p(\{\omega_{i},Z_{i}\},\{1/\tau^{2}_{gk}\}\mid\{Y_{i}\},\bm{\theta}^{(t)})$ which factorizes into, respectively, a collection of Polya-Gamma (PG), categorical, and Inverse-Gaussian random variables. Their conditional distributions are shown below,
@@
$p(Z_{i}=k\mid Y_{i},\bm{\theta})$ $\propto p_{ik}^{Y_{i}}(1-p_{ik})^{1-Y_{i}}\pi_{ik};\quad p_{ik}=\frac{\exp(\psi_{ik})}{1+\exp(\psi_{ik})},$ (A3b)
$p(\omega_{i}\mid Z_{i}=k,\bm{\theta})$ $\sim\mathrm{PG}\left(1,\psi_{ik}\right),$ (A3c)
@@
In practice, we rely on a generalized EM algorithm where $Q_{\beta}$ is improved versus maximized for computational reasons; we do so using a conjugate gradient solver initialized at $\check{\bm{\beta}}^{(t)}$.
F.2.2 Updating Moderator Parameters
To update the moderator parameters $\bm{\phi}$, we use the second cycle of the AECM algorithm where only the $Z_{i}$ are treated as missing data. The $E$-step involves recomputing the group membership probabilities, i.e., $p(Z_{i}\mid Y_{i},\bm{\beta}^{(t+1)},\bm{\phi}^{(t)})$, given the updates in the first cycle. The implied $Q$-function is shown below,

### v5

@@ 5.3 Group Membership
The right panel of the figure shows the distribution of other respondent characteristics. In general, Group 2 in the two-group analysis and Group 3 in the three-group analysis consist of those who live in ZIP codes with few immigrants and have lower educational achievements. For the three-group analysis, those in Group 2 tend to be Republicans, whereas those in Group 1 are more likely to be Democrats. This is consistent with the finding of a larger penalty for entry without legal authorization in Group 2. Group 3 contains a mix of political ideologies, though it has more respondents who identify as Undecided/Independent/Other or not strong Republican than the other two groups.
Which respondent characteristics are predictive of the group membership? In addition to the covariate distribution for each group shown in Fig 5, we can also find how important each moderator is in predicting group membership, conditional on all other moderators. We examine how the predicted probabilities of group memberships change across respondents with different characteristics. Specifically, we estimate
$\mathbb{E}\left[\pi_{k}(X_{ij}=x_{1},\bm{X}_{i,-j})-\pi_{k}(X_{ij}=x_{0},\bm{X}_{i,-j})\right]$ (5.1)
where $x_{0}$ and $x_{1}$ are different values of covariate of interest $X_{ij}$. If $X_{ij}$ is a categorical variable, we set $x_{0}$ to the baseline level and $x_{1}$ to the level indicated on the vertical axis. If $X_{ij}$ is a continuous variable as in the case of the Hispanic prejudice score, then $x_{0}$ and $x_{1}$ represent the 25th and 75th percentile values. The solid arrows represent whether the corresponding 95% Bayesian credible interval covers zero or not. Section L of the Supplementary Material shows the effect of changing a moderator on the absolute value of the changes in predicted probabilities of group membership. In some cases, changing a moderator shows a small average change but a larger average of absolute changes.
Figure 6: The impact of moderator values on likelihood of being assigned to groups, for two-group (left two plots) and three-group (right three plots) analysis. Dark arrows indicate that there is a significant effect of the moderator on group membership, i.e., that the corresponding quantity defined in Equation (5.1) is statistically significant.
Consistent with the earlier findings, Fig 6 shows that those with high Hispanic prejudice scores tend to be part of Group 2 in the two-group analysis and Group 3 in the three-group analysis even after controlling for other moderators. These respondents are also less likely to be members of Group 1 in both analyses. Party ID also plays a statistically significant role (indicated by dark arrow). Controlling for other factors, in the three-group analysis, not strong Republicans tend to be part of Group 3 and more strong Democrats belonging to Group 1. On average, respondents in Group 1 have higher education in both analyses.
@@
All of the AMIE effects found are quite small, so we do not present those results here. According to the skill-premium theory of Newman and Malhotra (2019), we expect to find an interaction between job and country or education and country, in at least some groups. Unfortunately, our analysis does not find support for this hypothesis.
@@ 5.4 Comparison to an alternative method
$\mathrm{CAMCE}_{j}(l,l^{\prime};\bm{X}_{i})=\sum_{k=1}^{K}\delta_{jk}(l,l^{\prime})\pi_{k}(\bm{X}_{i})$ (5.2)
@@
Section B of the Supplementary Material compares the estimated CAMCE obtained from our method and Robinson and Duch (2024)’s (cjbart) using the same moderators and treatments. Our method discovers a considerable degree of heterogeneity in the CAMCEs whereas cjbart shows limited treatment effect variation for most countries. Under our model, the estimated heterogeneous effects are more strongly associated with predictors than cjbart; for example, our method finds a clear association, on average, between the estimated CAMCE and prejudice or party identification whereas cjbart does not.
@@ 6 Concluding Remarks
Figure A1: Comparison of discovered heterogeneous effects (the conditional average marginal component effects or CAMCEs) between the proposed method and the BART-based method cjbart. In both plots the y-axis corresponds to values estimated, either by our method (right) or by cjbart (Robinson and Duch, 2024) (left). The plots show the estimated effect of Iraq as compared to the baseline of Germany. In the top figure, the x-axis and color corresponds to the categories of individuals based on the quartile of their Hispanic prejudice score. In the bottom figure, the x-axis and color corresponds to party ID.
@@
Lemma C.1 (Deviance Decomposition for Exponential Family).
@@
Proposition D.1 (Finding maximally heterogeneous groups with moderators).
@@
Lemma D.1 (Entropy of Groups with Respect to Moderators).
@@
Supplementary Material E Inclusion of Higher Order Interactions
Here we illustrate how the model and regularization penalties in Section 3.3 can be extended to include higher order interactions in a straightforward manner. We show below the model including all higher order interactions, and including only a subset is direct.
Let $\mathcal{J}=\{1,\dots,J\}$ be the set of $J$ factors and let $\mathcal{T}$ be the set of all possible assignments on the $\mathcal{J}$ factors. Then our model for $\psi_{k}(\bm{T}_{i})$ with all interactions among factors is
$\qquad\ +\ \cdots+\sum_{\bm{t}\in\mathcal{T}}\mathbf{1}\{\bm{T}_{i}=\bm{t}\}\beta^{12\cdots K}_{k\bm{t}}$
$\ =\ \mu+\tilde{\bm{T}}_{i}^{\top}\bm{\beta}_{k}.$
In the above formulation, $\beta^{12\cdots K}_{k\bm{t}}$ is the $K$-way interaction coefficient in cluster $k$ for assignment $\bm{t}$.
Let $\mathcal{T}_{-j}$ be the set of all possible assignments on the $\mathcal{J}$ factors except for factor $j$. With some slight notation abuse by letting $\beta^{12\cdots K}_{kl\bm{t}_{-j}}$ be the $K$-way interaction coefficient in cluster $k$ for assignment $l$ for $j$ and $\bm{t}_{j}$ for the other $J-1$ factors, the ANOVA-type sum-to-zero constraints extend as follows:
$\sum_{l=0}^{L_{j}-1}\beta^{j}_{kl}=0,\sum_{l=0}^{L_{j}-1}\beta^{jj^{\prime}}_{kll^{\prime}}=\sum_{l^{\prime}=0}^{L_{j^{\prime}}-1}\beta^{jj^{\prime}}_{kll^{\prime}}=0,\dots,\sum_{l=0}^{L_{j}-1}\beta^{12\cdots K}_{kl\bm{t}_{-j}}=0$ (A1)
for $j,j^{\prime}=1,2,\ldots,J$ with $j^{\prime}>j$ and for all $\bm{t}_{-j}\in\mathcal{T}_{-j}$. We write them compactly as,
[…]
where each row of $\bm{C}^{\top}\bm{\beta}_{k}$ corresponds to one of the constraints given in Equation (A1).
For the structured sparsity, we have penalties of the form
$\sum_{j=1}^{J}\sum_{l_{j}=1}^{L_{j}}\sum_{l_{j}^{\prime}>l_{j}}^{L_{j}}\sqrt{(\beta^{j}_{l_{j}}-\beta^{j}_{l_{j}^{\prime}})^{2}+\sum_{j^{\prime}\neq j}\sum_{l_{j^{\prime}}=1}^{L_{j^{\prime}}}(\beta^{jj^{\prime}}_{l_{j}l_{j^{\prime}}}-\beta^{j{j^{\prime}}}_{l_{j}^{\prime}l_{j^{\prime}}})^{2}+\cdots+\sum_{\bm{t}_{-j}\in\mathcal{T}_{-j}}(\beta^{12\cdots K}_{l_{j}\bm{t}_{-j}}-\beta^{12\cdots K}_{l_{j}^{\prime}\bm{t}_{-j}})^{2}}$
This will have $\sum_{j=1}^{J}L_{j}(L_{j}-1)/2$ terms, $L_{j}(L_{j}-1)/2$ terms for the $j$th factor.
For illustration, consider a simple example with one group and three factors—factor one has three levels, factor two has two levels, and factor three has two levels. In this case, our penalty contains 5 terms,
$\begin{split}&\sum_{l_{1}=1}^{L_{1}}\sum_{l_{1}^{\prime}>l_{1}}^{L_{1}}\sqrt{(\beta^{1}_{l_{1}}-\beta^{1}_{l_{1}^{\prime}})^{2}+\sum_{l_{2}=1}^{L_{2}}(\beta^{12}_{l_{1}l_{2}}-\beta^{12}_{l_{1}^{\prime}l_{2}})^{2}+\sum_{l_{3}=1}^{L_{3}}(\beta^{13}_{l_{1}l_{3}}-\beta^{13}_{l_{1}^{\prime}l_{3}})^{2}+\sum_{l_{2}=1}^{L_{2}}\sum_{l_{3}=1}^{L_{3}}(\beta^{123}_{l_{1}l_{2}l_{3}}-\beta^{123}_{l_{1}^{\prime}l_{2}l_{3}})^{2}}\\ ~{}+~{}&\sum_{l_{2}=1}^{L_{2}}\sum_{l_{2}^{\prime}>l_{2}}^{L_{2}}\sqrt{(\beta^{1}_{l_{2}}-\beta^{2}_{l_{2}^{\prime}})^{2}+\sum_{l_{2}=1}^{L_{2}}(\beta^{12}_{l_{1}l_{2}}-\beta^{12}_{l_{1}l_{2}^{\prime}})^{2}+\sum_{l_{3}=1}^{L_{3}}(\beta^{23}_{l_{2}l_{3}}-\beta^{23}_{l_{2}^{\prime}l_{3}})^{2}+\sum_{l_{1}=1}^{L_{1}}\sum_{l_{3}=1}^{L_{3}}(\beta^{123}_{l_{1}l_{2}l_{3}}-\beta^{123}_{l_{1}l_{2}^{\prime}l_{3}})^{2}}\\ \\ ~{}+~{}&\sum_{l_{3}=1}^{L_{3}}\sum_{l_{3}^{\prime}>l_{3}}^{L_{3}}\sqrt{(\beta^{1}_{l_{3}}-\beta^{3}_{l_{3}^{\prime}})^{2}+\sum_{l_{3}=1}^{L_{2}}(\beta^{13}_{l_{1}l_{3}}-\beta^{12}_{l_{1}l_{3}^{\prime}})^{2}+\sum_{l_{1}=1}^{L_{3}}(\beta^{13}_{l_{1}l_{3}}-\beta^{23}_{l_{1}l_{3}^{\prime}})^{2}+\sum_{l_{1}=1}^{L_{1}}\sum_{l_{2}=1}^{L_{2}}(\beta^{123}_{l_{1}l_{2}l_{3}}-\beta^{123}_{l_{1}l_{2}l_{3}^{\prime}})^{2}}\\ \end{split}$
The first three terms encourages the pairwise fusion of the levels of factor one whereas the fourth encourages the fusion of the two levels of factor two and the fifth encourages the fusion of the two levels of factor three.
Using the sum of Euclidean norms of quadratic forms, we can write the penalty as
$||\bm{\beta}^{\top}\bm{F}_{1}\bm{\beta}||_{2}+\ ||\bm{\beta}^{\top}\bm{F}_{2}\bm{\beta}||_{2}+\ ||\bm{\beta}^{\top}\bm{F}_{3}\bm{\beta}||_{2}+\ ||\bm{\beta}^{\top}\bm{F}_{4}\bm{\beta}||_{2}+\ ||\bm{\beta}^{\top}\bm{F}_{5}\bm{\beta}||_{2},$
[…]
More generally, for a fully interacted model we will have $\sum_{j=1}^{J}L_{j}(L_{j}-1)/2=G$ terms,
$\sum_{g=1}^{G}||\bm{\beta}^{\top}\bm{F}_{g}\bm{\beta}||_{2}.$
Supplementary Material F Propriety of the Structured Sparse Prior
@@
By applying Theorem 1 and noting that the null spaces of $\bm{A}^{T}\bm{A}$ and $\bm{A}$ are identical, the integral of $\tilde{k}(\tilde{\bm{\beta}})$ is finite if and only if $\bm{Q}\mathcal{B}_{\bm{C}^{\top}}$ is full column rank, where $\bm{Q}^{\top}=[\bm{Q}^{\top}_{1},\cdots,\bm{Q}^{\top}_{G}]$. We demonstrate this fact in two steps. First, there exists a permutation matrix $\bm{P}_{Q}$ such that $\bm{P}_{Q}\bm{Q}$ has a block diagonal structure with $J+1$ diagonal blocks. The first $J$ blocks corresponding to the main terms for each factor $j$ and the last block corresponds to all interaction terms. The null space of each block is spanned by the vector $\bm{1}$ as the corresponding block of $\bm{P}_{Q}\bm{Q}$ is a (transposed) orientated incidence matrix of a fully connected graph. Thus, the null space of $\bm{P}_{Q}\bm{Q}$, and hence $\bm{Q}$, is spanned by the $J+1$ columns of a block diagonal matrix with $\bm{1}$ on each block. Second, consider the linear constraints $\bm{C}^{\top}\bm{\beta}=\bm{0}$. The only vector to satisfy this constraint and lie in the null space of $\bm{Q}$ must be $\bm{0}$ as, for each block, the only vector proportional to $\bm{1}$ and satisfying the corresponding sum-to-zero constraints must be $\bm{0}$. Thus, $\bm{Q}\mathcal{B}_{\bm{C}^{\top}}$ is full column rank and the prior is proper. ∎
Supplementary Material G Derivations for the Basic Model
@@
G.1 Removing the Linear Constraints
@@
Given the similarity of the unconstrained and constrained problems and for notational simplicity, we present all results herein dropping the second “tilde” notation on $\tilde{\bm{T}}_{i}$ and the “tilde” on $\bm{\beta}_{k}$ and note that, once estimated, $\tilde{\bm{\beta}}_{k}$ is projected back into the original space for the reported coefficients, average marginal component effects, etc. The results of Appendix J on approximating $\tilde{\bm{\beta}}_{k}$ as multivariate Gaussian imply that $\bm{\beta}_{k}$ will have a (singular) multivariate Gaussian distribution.
G.2 Expectation Maximization Algorithm
This section considers inference after removing the linear constraints as discussed in the prior subsection. Algorithm A1 summarizes our approach to maximizing Equation (3.10). Each iteration of our AECM algorithm involves two cycles where the data augmentation scheme enables iterative updating of the treatment effect parameters $\bm{\beta}$ and moderators $\bm{\phi}$. $\bm{\theta}$ collects both sets of parameters.
@@ G.2 Expectation Maximization Algorithm
Initialize Parameters: $\bm{\theta}^{(0)}$, i.e. $\bm{\beta}^{(0)}$ and $\bm{\phi}^{(0)}$; Appendix G.5 provides details.
@@
     1a. $E$-Step: Find the conditional distributions of $\{Z_{i},\omega_{i}\}_{i=1}^{N}$ and $\{\{\tau^{2}_{gk}\}_{g=1}^{G}\}_{k=1}^{K}$ given $\{Y_{i},\bm{X}_{i},\bm{T}_{i}\}$ and $\bm{\theta}^{(t)}$ (Eq. (A1)). Derive $Q_{\beta}(\bm{\beta},\bm{\theta}^{(t)})$ (Eq. (G.2.1)).
@@
     2a. $E$-Step: Find $p(Z_{i}=k\mid Y_{i},\bm{X}_{i},\bm{T}_{i},\bm{\beta}^{(t+1)},\bm{\phi}^{(t)})$. Derive $Q_{\phi}(\bm{\phi},\{\bm{\beta}^{(t+1)},\bm{\phi}^{(t)}\})$ (Eq. (G.2.2)).
@@
     3. Stop if $\log p\left(\bm{\theta}^{(t+1)}|\{Y_{i},\bm{X}_{i},\bm{T}_{i}\}_{i=1}^{N}\right)-\log p\left(\bm{\theta}^{(t)}|\{Y_{i},\bm{X}_{i},\bm{T}_{i}\}_{i=1}^{N}\right)<\epsilon_{1}$ (Eq. (3.10)) or
$||\bm{\theta}^{(t+1)}-\bm{\theta}^{(t)}||_{\infty}<\epsilon_{2}$.
G.2.1 Updating Treatment Effect Parameters
@@
$p(Y_{i},\omega_{i}\mid Z_{i},\bm{X}_{i},\bm{T}_{i})\ \propto\ \frac{1}{2}\exp\left\{\left(Y_{i}-\frac{1}{2}\right)\psi_{Z_{i}}(\bm{T}_{i})-\frac{\omega_{i}}{2}\left[\psi_{Z_{i}}(\bm{T}_{i})\right]^{2}\right\}f_{PG}(\omega_{i}\mid 1,0),$ (A1a)
@@
where all expectations are taken over the conditional distribution of the missing data given the current parameter estimates. We note that the $E$-Step involves computing $p(\{\omega_{i},Z_{i}\},\{1/\tau^{2}_{gk}\}\mid\{Y_{i},\bm{X}_{i},\bm{T}_{i}\},\bm{\theta}^{(t)})$ which factorizes into, respectively, a collection of Polya-Gamma (PG), categorical, and Inverse-Gaussian random variables. Their conditional distributions are shown below,
@@
$p(Z_{i}=k\mid Y_{i},\bm{X}_{i},\bm{T}_{i},\bm{\theta})$ $\propto p_{ik}^{Y_{i}}(1-p_{ik})^{1-Y_{i}}\pi_{ik};\quad p_{ik}=\frac{\exp(\psi_{ik})}{1+\exp(\psi_{ik})},$ (A3b)
$p(\omega_{i}\mid Z_{i}=k,\bm{X}_{i},\bm{T}_{i},\bm{\theta})$ $\sim\mathrm{PG}\left(1,\psi_{ik}\right),$ (A3c)
@@
One could reply on a generalized EM algorithm where $Q_{\beta}$ is improved versus maximized for computational reasons, e.g. by using a conjugate gradient solver initialized at $\check{\bm{\beta}}^{(t)}$.
G.2.2 Updating Moderator Parameters
To update the moderator parameters $\bm{\phi}$, we use the second cycle of the AECM algorithm where only the $Z_{i}$ are treated as missing data. The $E$-step involves recomputing the group membership probabilities, i.e., $p(Z_{i}\mid Y_{i},\bm{X}_{i},\bm{T}_{i},\bm{\beta}^{(t+1)},\bm{\phi}^{(t)})$, given the updates in the first cycle. The implied $Q$-function is shown below,